OpenStudy (brittany16):

Which operation would isolate the variable in this equation? 6m = -36

OpenStudy (anonymous):

Division

jimthompson5910 (jim_thompson5910):

You would divide both sides by 6 to get 6m = -36 6m/6 = -36/6 m = -6

OpenStudy (anonymous):

divide both sides by 6